# Adequately Tailoring Age Verification Regulations

### v1

@@ Abstract.
In response to growing concerns about harms to children online, 25 states in the U.S. have enacted laws requiring age verification to restrict minors access to online pornography. In June 2025, the Supreme Court decision in Free Speech Coalition v. Paxton upheld the constitutionality of Texas H.B. 1181, one of the most constitutionally vulnerable of these age verification laws, holding that it was subject to and satisfied intermediate scrutiny and the requirement that age verification regulations be “adequately tailored”. However, the decision leaves unresolved practical challenges. What is the current state of age verification legislation in the United States? How can “adequate tailoring” be interpreted in a way that is accessible to non-legal experts, particularly those in technical and engineering domains? What age verification approaches are used today, what infrastructures and standards support them, and what tradeoffs do they introduce? This paper addresses those questions by proposing an analytical model to interpret “adequate tailoring” from multiple perspectives with associated governmental goals and interests, and by applying that model to evaluate both current state laws and widely used verification methods. This paper’s major contributions include: (1) we mapped the current U.S. age-verification legislative landscape; (2) we introduce an analytical model to analyze “adequate tailoring” for age verification and potential application to other online regulatory policies; and (3) we analyze the main technical approaches to age verification, highlighting the practical challenges and tradeoffs from a technical perspective. Further, while we focus on U.S. State laws, the principles underlying our framework are applicable to age-verification debates and methods worldwide.
@@ 1. Introduction
  • RQ2: How should “adequate tailoring” be interpreted? To address this question, we identify key governmental goals, and identify tensions among four properties of age assurance systems: Assurance, Business Convenience, Consumer Convenience, and Data Protection together with related governmental interests. (Section 3)
@@
Recent state legislation reflects a rapidly consolidating approach to age verification, with a clear emphasis on enforcing minimum age thresholds—most commonly 18 years—through structured verification mechanisms. Table 1 summarizes state statutes that specify age verification requirements and related privacy provisions within those statutes. States nominally permit a variety of verification approaches, and most states (eighteen) permit use of any “reasonable commercial method” to verify age. However, many states instead (or additionally) mention specific methods. In practice, excluding the general “reasonable commercial method” language, the regulatory landscape is dominated by ID-based methods. Of the twenty-five states with active age verification laws, sixteen states, including Arizona and South Dakota accept submission of traditional IDs (2112, 2025; 3424, 2024b), while thirteen states like Arkansas, Georgia, Indiana accept digital IDs (66, 2023; 351, 2024; 17, 2024). By contrast, specific mentions of non-ID-based approaches within the statute itself are far less common. Only three states, Nebraska, South Dakota, and Wyoming accept financial documents, while 18 states like Virginia accept commercial system based on transactional data or other databases (1515, 2023), and nine states like Florida accept other methods (3, 2024).
Almost all of the statues also address privacy and data retention (or non-retention) requirements. All states except Virginia (1515, 2023) have at least some form of privacy requirement. Eighteen states such as Alabama, Arkansas, and Florida, prohibit the retention of personal information after an age verification check has been completed (164, 2024; 66, 2023; 3, 2024). Five states including Arizona, Indiana, Missouri, South Dakota, and Texas prohibit retention of verification data at any time (2112, 2025; 66, 2023; 60-18, 2025; 3424, 2024b; 1181, 2023).
@@ 2.2. Protecting children from harmful Internet content: From Reno to Paxton
Although proof of age has been required for activities such as voting and marriage (Cross, 2004) for many years, the legislative emphasis on protecting children (from harmful content or from other harms) arose much later (Marsden, 2023).
Modern debates on age verification in the U.S. trace back to the Communications Decency Act of 1996 and the landmark Supreme Court decision in Reno v. American Civil Liberties Union (Marsden, 2023; Reno v. Am. C.L. Union, 521 U.S. 844, 882, 1997). The First Amendment of the US Constitution protects individuals from government restrictions on expression, and courts establish three levels of scrutiny (i.e., “strict scrutiny”, “intermediate scrutiny” and “rational basis review”) when reviewing laws that limit protected speech. In Reno, the Court struck down provisions restricting knowing “transmission” of“indecent” or “patently offensive” online speech to minors, holding that they imposed unconstitutional burdens on free expression for adults and failed in passing “strict scrutiny” (Reno v. Am. C.L. Union, 521 U.S. 844, 882, 1997). Since then, the tension between protecting children and safeguarding First Amendment rights has remained a contentious issue in both legislative and judicial arenas (Thierer, 2007). Apthorpe, Frischmann, and Shvartzshnaider examine the state of age gating before FSC v. Paxton (Frischmann et al., 2024).
@@ 3. A framework for interpreting “adequate tailoring”
However, how should websites, age verifiers, judges, and the public interpret “adequate tailoring”? especially given all those different governmental interests? In this section, we propose the analysis model to provide a structured way to reason about “adequate tailoring”. We first identify the government’s Goals/Interests by reading the statutes and surrounding contextual information (Section 3.1). Then, we identify properties of age verification systems that interact with those goals. In Section 3.2 We identify four properties – Assurance, Data Protection, Business Convenience, and Consumer Convenience – which we will later use to understand what is being optimized, what is being sacrificed, and how alternative approaches compare.
@@ 3.1. Categories of government goals
  • Privacy and Security Goal denotes the governmental interest in preserving individual privacy and safeguarding data security while pursuing. For instance, most age verification laws impose requirements to avoid retaining personal information after an age check or to prohibit ongoing data storage (see Table 1 “Privacy Requirements”).
  • Consumer Rights Protection Goal denotes the protection of lawful consumer rights in the pursuit of Target Goal. For instance, both Oklahoma and Kansas mentions “consumer protection” in the age verification bills (1959, 2024; 394, 2024). Generally, in the age-verification context, Consumer Rights Protection Goal includes safeguarding the rights of adults to access online pornography and other legally permissible materials.
@@ 4.1.1. Direct ID document upload
One of the most straightforward age verification methods is the direct upload of an official ID document. In this approach, users are prompted to provide a scan or photograph of a government-issued identification (e.g. driver’s license, passport) via a website or app. The system’s infrastructure typically includes an image capture interface and a document verification service, which technically will extract data from the ID using optical character recognition (OCR) or barcode scanning (Ligon, 2025). For example, U.S. driver’s licenses encode personal data in a PDF417 barcode per AAMVA standards, which can be read to obtain the date of birth and other details (of Motor Vehicle Administrators, 2025).
@@
Despite its effectiveness at directly checking an official birth date, the direct ID upload method comes with notable privacy and security concerns. On the privacy side, uploading a government ID exposes personal data well beyond just age (name, address, ID number, etc.), raising the stakes for data handling. Regarding assurance, service providers typically authenticate an uploaded ID copy through a pipeline that includes an automated check, human review, or both. At submission, systems run image-quality and tamper checks (e.g., detection of blurring, irregularities in fonts/spacing/edges, and evidence of digital editing), followed by template-based validation against known ID layouts and security features for the claimed issuing jurisdiction (e.g., expected placement of fields, UV/hologram regions—where visible—or microprint proxies) (Boult, 2025). Many vendors then parse the machine-readable zone: for U.S. driver’s licenses this commonly means decoding the AAMVA PDF417 barcode to cross-check the barcode payload (date of birth, name, ID number, address) against the human-readable fields and to flag internal inconsistencies that are common in forgeries (of Motor Vehicle Administrators, 2025). Where available, higher-assurance implementations also seek issuer-backed validation such as cryptographic signatures previously signed by the ID issuance authorities. In practice, however, uploaded physical license images often cannot be cryptographically verified end-to-end: many states’ barcodes do not contain the optional DMV-signed cryptographic digital signature. As a result, human review is frequently used as a backstop for borderline cases (e.g., low-confidence authenticity scores or suspected synthetic images), but manual inspection is inherently variable and can be defeated by high-quality counterfeits or well-crafted digital manipulations.
@@ 4.1.2. Verifiable digital credentials
A Digital ID is a long-term digital representation of an ID. One key potential benefit of digital IDs, as described in the ISO/IEC 18013 standard, are their ability to perform selective disclosure, which enables user to pick and choose subset of attributes information to share with relying parties, rather than sharing all ID information (International Organization for Standardization, 2017, 2021). The Notary provides an attestation of this property. For example, someone could reveal their birthdate, signed by the notary application, to prove their age. The Assurance on the ID’s validity, and the Data Protection of the system, all relies on the Notary: The relying party must trust an attestation about the ID made by the notary, and the the User must trust the Notary to handle their ID information privately and securely. We note that the wallet application where the credential is stored need not be the same location as where the ID is checked (see Figure 1). Implementations of this approach include Apple Wallet (Apple Inc., 2025), Dock Wallet (Dock Labs, 2024), Microsoft Entra, (Microsoft, 2025), or a Digital ID application offered by a U.S. State (e.g., (State of Colorado, 2025) or (of Motor Vehicles, [n. d.])).
@@
For either implementation of the technical Relative to direct ID upload, verifiable ID credentials can materially reduce data exposure by returning only the minimal assertion needed for the transaction, compared to transmitting a full ID. Selective disclosure limits the claims revealed, while ZKPs can further limit disclosure by proving only compliance with a threshold and potentially reducing linkability risks across verifications. Further, verifiable ID credentials support additional features including anti-replay nonces and optional revocation/status checks (Sporny et al., 2025), which further improve security controls and privacy safeguarding. However, these approaches introduce new accessibility and operation considerations. First, users may need compatible smartphones, wallets, and connectivity (depending on whether the flow supports offline verification), and recovery processes for lost devices can be complex. Some users (including those without qualifying state IDs, those with limited digital literacy, or those lacking modern devices) may be disadvantaged, suggesting a continued need for fallback mechanisms (e.g., in-person verification or alternative non-ID methods) to avoid exclusion. Second, from an implementer perspective, interoperability remains a practical challenge: relying parties must integrate wallet/presentation protocols and align with issuer trust frameworks, which can slow deployment compared to “upload an ID image” solutions. Third, usability remains a major challenge. Verifiable digital credentials rely on complex cryptographic mechanisms that can be difficult for non-expert users to understand, which may hinder broader adoption and dilute the practical benefits of their privacy-preserving features.
@@ 4.1.3. Facial age estimation
Facial age estimation offers a low-friction alternative to ID-based checks and are likely more convenient to consumers than ID-based alternatives. However, they present several material drawbacks that affect both assurance and privacy. First, age inference is inherently difficult, and report shows that age estimation accuracy is “strongly influenced by algorithm, sex, image quality, region-of-birth, age itself, and interactions between those factor” (Hanaoka et al., 2024). Variations in image quality, makeup, facial expression, and environmental factors such as lighting and camera angle can materially affect model performance (Haugg et al., 2025), and potentially resulting in false rejections of legitimate adults or false acceptances of older-looking minors. Second, accessibility can be a barrier. Users with certain disabilities or medical conditions may struggle with camera-based prompts such as maintaining a steady pose or aligning their face, and individuals with atypical facial features, facial paralysis, or assistive devices may face higher age estimation failure rates (James, 2024). Finally, even when vendors claim minimal retention, facial age estimation raises privacy concerns because it processes biometric data such as facial images and derived signals that are widely treated as highly sensitive.
Some facial age estimation deployments aditionally include liveness detection, which helps ensure the camera is capturing a real, present user rather than a photo or video replay. Liveness detection can be implemented either by prompting an action (having the user turn their head) or by inferring other signals from the capture (e.g. whether the camera is moving in small ways indicating the camera is held) (Alvargonzález, 2025). Action prompting is considered to have stronger assurance, but lower user convenience (Alvargonzález, 2025).
@@ 4.3. Are credentials device-binding or people-binding?
  • Facial age estimation. Facial age estimation is intended to be people-bound at the point of decision because it evaluates the live user in front of the camera and can incorporates liveness checks to reduce replay. However, an underage person might simply point the camera at an overage person and cause the “binding” to misfire on the wrong person.
@@
These requirements, particularly photo matching and periodic refresh, more directly advance the Act’s primary target goal of “protecting children from social media” as suggested by the Act title. Photo matching introduces an explicit people-binding element that can increase assurance by reducing circumvention via shared devices, borrowed accounts, or reused ID images. The 60-minute session limit further constrains reuse of a single verification event and helps mitigate ongoing access by minors after an adult has verified. Although the fake ID issue discussed in section 4.1.1 may still be possible, these design choices still strengthen effective prevention of minor from accessing targeted websites by raising the Assurance level of age gating. However, such achievements come with significant decreases in Data Protection, Consumer Convenience, and Business Convenience, and finally impede Privacy and Security Goal and other Residual Interests.
@@
Further, the 60-minute verification session further force users to re-verify their identity repeatedly, disrupting the online experience and creating excessive inconvenience. Rather than enhancing the effectiveness of age-verification regimes, such rules introduce needless friction that can discourage compliance. In sum, when verification regimes demand repeated disclosures of sensitive data, impose short session limits, or require intrusive biometric checks, they substantially increase the burden on ordinary users. These measures reduce Consumer Convenience by introducing friction, privacy risks, and technical obstacles that make lawful access more difficult.
Finally, Tennessee’s requirement to retain seven years of “anonymized age-verification data” may materially lower Business Convenience by imposing ongoing compliance, storage, and audit burdens that extend well beyond the real-time act of age gating. First, the definition of “anonymized age verification data” and “personally identifying information” are unclear or inaccurate. For instance, depending on the technical method used, supposedly “anonymized” attributes such as date of birth ranges, device identifiers, or partial government ID numbers may still be re-identifiable when combined with other datasets (Rocher et al., 2019), especially over a 7-year window. This definitional ambiguity risks leaving large volumes of quasi-identifiable information vulnerable to breach or misuse. Second, 7 years is a long period for storing any form of sensitive or quasi-sensitive data, especially given the accelerating pace of data breaches and cyberattacks. The longer data is retained, the more likely it is to be exposed through hacking, insider misuse, or even changes in legal regimes that alter the permissible uses of the data. In spite of these privacy concerns, maintaining long-horizon records at scale requires additional logging infrastructure, secure storage and backup, access controls, retention enforcement, and internal governance processes to ensure data remains available, accurate, and defensible over time. These requirements can be especially costly for smaller publishers and for sites that rely on third-party verification vendors, because contractual arrangements, data pipelines, and audit responsibilities must be structured to preserve the mandated records while still complying with minimization obligations. In practice, the combination of long retention and evidentiary expectations increases fixed compliance costs, raises vendor-management complexity, and heightens legal exposure, thereby reducing Business Convenience.

### v2

@@ Abstract.
In response to growing concerns about the harms to children online, especially regarding access to online pornography, a new wave of regulatory action has swept across the United States. To date, 25 states have implemented laws aimed at restricting minors’ access to such content by requiring websites to implement “reasonable age verification.” In June 2025, the Supreme Court’s decision in Free Speech Coalition v. Paxton upheld the constitutionality of Texas H.B. 1181 – one of the most constitutionally vulnerable of these age verification laws – and found that it was subject to (and passed the standard of) intermediate scrutiny, with a need to “adequately tailor” age verification regulations. Yet, Free Speech Coalition v. Paxton leaves unresolved practical challenges posed by state-level age-verification regimes.
What is the state of age verification legislation in the US? How can the constitutional concept of “adequate tailoring” be interpreted in a way that is accessible and meaningful for non-legal experts, particularly those in technical and engineering domains who are attempting to interpret these laws in the context of technical trade-offs? What age-verification approaches are widely used today, what technical infrastructures and standards do they rely on, and how do they contribute to these concerns? This paper addresses those questions by proposing an analytical model to interpret “adequate tailoring” from multiple perspectives with associated governmental goals and interests, and properties of age verification systems, and by applying that model to evaluate both current state laws and widely used verification methods.
This paper’s major contributions include: (1) we map the current U.S. age-verification legislative landscape; (2) we introduce an analytical model to analyze “adequate tailoring” for age verification and potential application to other online regulatory policies; and (3) we analyze the main technical approaches to age verification, highlighting the practical challenges and tradeoffs from a technical perspective. Further, while we focus on U.S. state laws, the principles underlying our framework are applicable to age-verification debates and methods worldwide.
@@ 1. Introduction
  • RQ2: How should “adequate tailoring” be interpreted? To address this question, we identify key governmental goals and tensions among four properties of age assurance systems: Assurance, Business Convenience, Consumer Convenience, and Data Protection together with related governmental interests. (Section 3)
@@
Recent state legislation reflects a rapidly consolidating approach to age verification, with a clear emphasis on enforcing minimum age thresholds—most commonly 18 years—through structured verification mechanisms. Table 1 summarizes state statutes that specify age verification requirements and related privacy provisions within those statutes. States nominally permit a variety of verification approaches, and most states (eighteen) permit use of any “reasonable commercial method” to verify age. However, many states instead (or additionally) mention specific methods. In practice, excluding the general “reasonable commercial method” language, the regulatory landscape is dominated by ID-based methods. Of the twenty-five states with active age verification laws, sixteen states, including Arizona and South Dakota accept submission of traditional IDs (2112, 2025; 3424, 2024b), while thirteen states like Arkansas, Georgia, Indiana accept digital IDs (66, 2023; 351, 2024; 17, 2024). By contrast, specific mentions of non-ID-based approaches within the statute itself are far less common. Only three states, Nebraska, South Dakota, and Wyoming accept financial documents, while 18 states like Virginia accept commercial systems based on transactional data or other databases (1515, 2023), and nine states like Florida accept other methods (3, 2024).
Almost all of the statutes also address privacy and data retention (or non-retention) requirements. All states except Virginia (1515, 2023) have at least some form of privacy requirement. Eighteen states such as Alabama, Arkansas, and Florida, prohibit the retention of personal information after an age verification check has been completed (164, 2024; 66, 2023; 3, 2024). Five states including Arizona, Indiana, Missouri, South Dakota, and Texas prohibit retention of verification data at any time (2112, 2025; 66, 2023; 60-18, 2025; 3424, 2024b; 1181, 2023).
@@ 2.2. Protecting Children from Harmful Internet Content: From Reno to Paxton
Although proof of age has been required for activities such as voting and marriage (Cross, 2004) for many years, the legislative emphasis on protecting children (from harmful content or from other harms) arose much later (Marsden, 2023). Modern debates on age verification in the U.S. trace back to the Communications Decency Act of 1996 and the landmark Supreme Court decision in Reno v. American Civil Liberties Union (Marsden, 2023; Reno v. Am. C.L. Union, 521 U.S. 844, 882, 1997). The First Amendment of the US Constitution protects individuals from government restrictions on expression, and courts establish three levels of scrutiny (i.e., “strict scrutiny”, “intermediate scrutiny” and “rational basis review”) when reviewing laws that limit protected speech. In Reno, the Court struck down provisions restricting knowing “transmission” of “indecent” or “patently offensive” online speech to minors, holding that they imposed unconstitutional burdens on free expression for adults and failed in passing “strict scrutiny” (Reno v. Am. C.L. Union, 521 U.S. 844, 882, 1997). Since then, the tension between protecting children and safeguarding First Amendment rights has remained a contentious issue in both legislative and judicial arenas (Thierer, 2007). Apthorpe, Frischmann, and Shvartzshnaider examine the state of age gating before FSC v. Paxton (Frischmann et al., 2024).
@@ 3. A Framework for Interpreting “Adequate Tailoring”
However, how should websites, age verifiers, judges, and the public interpret “adequate tailoring”, especially given all those different governmental interests? In this section, we propose the analysis model to provide a structured way to reason about “adequate tailoring”. We first identify the government’s Goals/Interests by reading the statutes and surrounding contextual information (Section 3.1). Then, we identify properties of age verification systems that interact with those goals. In Section 3.2 we identify four properties – Assurance, Data Protection, Business Convenience, and Consumer Convenience – which we will later use to understand what is being optimized, what is being sacrificed, and how alternative approaches compare.
@@ 3.1. Categories of Government Goals
  • Privacy and Security Goal denotes the governmental interest in preserving individual privacy and safeguarding data security while pursuing other objectives. For instance, most age verification laws impose requirements to avoid retaining personal information after an age check or to prohibit ongoing data storage (see Table 1 “Privacy Requirements”).
  • Consumer Rights Protection Goal denotes the protection of lawful consumer rights. For instance, both Oklahoma and Kansas mention “consumer protection” in the age verification bills (1959, 2024; 394, 2024). Generally, in the age-verification context, Consumer Rights Protection Goal includes safeguarding the rights of adults to access online pornography and other legally permissible materials.
@@ 4.1.1. Direct ID document upload
One of the most straightforward age verification methods is the direct upload of an official ID document. In this approach, users are prompted to provide a scan or photograph of a government-issued identification (e.g. driver’s license, passport) via a website or app. The system’s infrastructure typically includes an image capture interface and a document verification service, which will technically extract data from the ID using optical character recognition (OCR) or barcode scanning (Ligon, 2025). For example, U.S. driver’s licenses encode personal data in a PDF417 barcode per AAMVA standards, which can be read to obtain the date of birth and other details (of Motor Vehicle Administrators, 2025).
@@
Despite its effectiveness at directly checking an official birth date, the direct ID upload method comes with notable privacy and security concerns. On the privacy side, uploading a government ID exposes personal data well beyond just age (name, address, ID number, etc.), raising the stakes for data handling. Regarding assurance, service providers typically authenticate an uploaded ID copy through a pipeline that includes an automated check, human reviews, or both. At submission, systems run image-quality and tamper checks (e.g., detection of blurring, irregularities in fonts/spacing/edges, and evidence of digital editing), followed by template-based validation against known ID layouts and security features for the claimed issuing jurisdiction (e.g., expected placement of fields, UV/hologram regions—where visible—or microprint proxies) (Boult, 2025). Many vendors then parse the machine-readable zone: for U.S. driver’s licenses this commonly means decoding the AAMVA PDF417 barcode to cross-check the barcode payload (date of birth, name, ID number, address) against the human-readable fields and to flag internal inconsistencies that are common in forgeries (of Motor Vehicle Administrators, 2025). Where available, higher-assurance implementations also seek issuer-backed validation such as cryptographic signatures previously signed by the ID issuance authorities. In practice, however, uploaded physical license images often cannot be cryptographically verified end-to-end: many states’ barcodes do not contain the optional DMV-signed cryptographic digital signature. As a result, human review is frequently used as a backstop for borderline cases (e.g., low-confidence authenticity scores or suspected synthetic images), but manual inspection is inherently variable and can be defeated by high-quality counterfeits or well-crafted digital manipulations.
@@ 4.1.2. Verifiable digital credentials
A Digital ID is a long-term digital representation of an ID. One key potential benefit of digital IDs, as described in the ISO/IEC 18013 standard, is their ability to perform selective disclosure, which enables users to pick and choose subset of attributes information to share with relying parties, rather than sharing all ID information (International Organization for Standardization, 2017, 2021). The Notary provides an attestation of this property. For example, someone could reveal their birthdate, signed by the notary application, to prove their age. The Assurance on the ID’s validity, and the Data Protection of the system, all rely on the Notary: The relying party must trust an attestation about the ID made by the notary, and the User must trust the Notary to handle their ID information privately and securely. We note that the wallet application where the credential is stored needs not be the same location as where the ID is checked (see Figure 1). Implementations of this approach include Apple Wallet (Apple Inc., 2025), Dock Wallet (Dock Labs, 2024), Microsoft Entra, (Microsoft, 2025), or a Digital ID application offered by a U.S. State (e.g., (State of Colorado, 2025) or (of Motor Vehicles, [n. d.])).
@@
Relative to direct ID upload, verifiable ID credentials can materially reduce data exposure by returning only the minimal assertion needed for the transaction, compared to transmitting a full ID. Selective disclosure limits the claims revealed, while ZKPs can further limit disclosure by proving only compliance with a threshold and potentially reducing linkability risks across verifications. Further, verifiable ID credentials support additional features including anti-replay nonces and optional revocation/status checks (Sporny et al., 2025), which further improve security controls and privacy safeguarding. However, these approaches introduce new accessibility and operation considerations. First, users may need compatible smartphones, wallets, and connectivity (depending on whether the flow supports offline verification), and recovery processes for lost devices can be complex. Some users (including those without qualifying state IDs, those with limited digital literacy, or those lacking modern devices) may be disadvantaged, suggesting a continued need for fallback mechanisms (e.g., in-person verification or alternative non-ID methods) to avoid exclusion. Second, from an implementer perspective, interoperability remains a practical challenge: relying parties must integrate wallet/presentation protocols and align with issuer trust frameworks, which can slow deployment compared to “upload an ID image” solutions. Third, usability remains a major challenge. Verifiable digital credentials rely on complex cryptographic mechanisms that can be difficult for non-expert users to understand, which may hinder broader adoption and dilute the practical benefits of their privacy-preserving features.
@@ 4.1.3. Facial age estimation
Facial age estimation offers a low-friction alternative to ID-based checks and is likely more convenient to consumers than ID-based alternatives. However, they present several material drawbacks that affect both assurance and privacy. First, age inference is inherently difficult, and report shows that age estimation accuracy is “strongly influenced by algorithm, sex, image quality, region-of-birth, age itself, and interactions between those factors” (Hanaoka et al., 2024). Variations in image quality, makeup, facial expression, and environmental factors such as lighting and camera angle can materially affect model performance (Haugg et al., 2025), and potentially result in false rejections of legitimate adults or false acceptances of older-looking minors. Second, accessibility can be a barrier. Users with certain disabilities or medical conditions may struggle with camera-based prompts such as maintaining a steady pose or aligning their face, and individuals with atypical facial features, facial paralysis, or assistive devices may face higher age estimation failure rates (James, 2024). Finally, even when vendors claim minimal retention, facial age estimation raises privacy concerns because it processes biometric data such as facial images and derived signals that are widely treated as highly sensitive.
Some facial age estimation deployments additionally include liveness detection, which helps ensure the camera is capturing a real, present user rather than a photo or video replay. Liveness detection can be implemented either by prompting an action (having the user turn their head) or by inferring other signals from the capture (e.g. whether the camera is moving in small ways indicating the camera is held) (Alvargonzález, 2025). Action prompting is considered to have stronger assurance, but lower user convenience (Alvargonzález, 2025).
@@ 4.3. Are Credentials Device-Binding or People-Binding?
  • Facial age estimation. Facial age estimation is intended to be people-bound at the point of decision because it evaluates the live user in front of the camera and can incorporate liveness checks to reduce replay. However, an underage person might simply point the camera at an overage person and cause the “binding” to misfire on the wrong person.
@@
These requirements, particularly photo matching and periodic refresh, more directly advance the Act’s primary target goal of “protecting children from social media” as suggested by the Act title. Photo matching introduces an explicit people-binding element that can increase assurance by reducing circumvention via shared devices, borrowed accounts, or reused ID images. The 60-minute session limit further constrains reuse of a single verification event and helps mitigate ongoing access by minors after an adult has verified. Although the fake ID issue discussed in section 4.1.1 may still be possible, these design choices still strengthen effective prevention of minors from accessing targeted websites by raising the Assurance level of age gating. However, such achievements come with significant decreases in Data Protection, Consumer Convenience, and Business Convenience, and finally impede Privacy and Security Goal and other Residual Interests.
@@
Further, the 60-minute verification session further forces users to re-verify their identity repeatedly, disrupting the online experience and creating excessive inconvenience. Rather than enhancing the effectiveness of age-verification regimes, such rules introduce needless friction that can discourage compliance. In sum, when verification regimes demand repeated disclosures of sensitive data, impose short session limits, or require intrusive biometric checks, they substantially increase the burden on ordinary users. These measures reduce Consumer Convenience by introducing friction, privacy risks, and technical obstacles that make lawful access more difficult.
Finally, Tennessee’s requirement to retain seven years of “anonymized age-verification data” may materially lower Business Convenience by imposing ongoing compliance, storage, and audit burdens that extend well beyond the real-time act of age gating. First, the definition of “anonymized age verification data” and “personally identifying information” are unclear or inaccurate. For instance, depending on the technical method used, supposedly “anonymized” attributes such as date of birth ranges, device identifiers, or partial government ID numbers may still be re-identifiable when combined with other datasets (Rocher et al., 2019), especially over a 7-year window. This definitional ambiguity risks leaving large volumes of quasi-identifiable information vulnerable to breach or misuse. Second, 7 years is a long period for storing any form of sensitive or quasi-sensitive data, especially given the accelerating pace of data breaches and cyberattacks. The longer data are retained, the more likely it is to be exposed through hacking, insider misuse, or even changes in legal regimes that alter the permissible uses of the data. In spite of these privacy concerns, maintaining long-horizon records at scale requires additional logging infrastructure, secure storage and backup, access controls, retention enforcement, and internal governance processes to ensure data remain available, accurate, and defensible over time. These requirements can be especially costly for smaller publishers and for sites that rely on third-party verification vendors, because contractual arrangements, data pipelines, and audit responsibilities must be structured to preserve the mandated records while still complying with minimization obligations. In practice, the combination of long retention and evidentiary expectations increases fixed compliance costs, raises vendor-management complexity, and heightens legal exposure, thereby reducing Business Convenience.